### v1

# Instructing Large Language Models for Low-Resource Languages: A Systematic Study for Basque

Abstract
Instructing language models with user intent requires large instruction datasets, which are only available for a limited set of languages. In this paper, we explore alternatives to conventional instruction adaptation pipelines in low-resource scenarios. We assume a realistic scenario for low-resource languages, where only the following are available: corpora in the target language, existing open-weight multilingual base and instructed backbone LLMs, and synthetically generated instructions sampled from the instructed backbone. We present a comprehensive set of experiments for Basque that systematically study different combinations of these components evaluated on benchmarks and human preferences from participants. Our conclusions show that target language corpora are essential, with synthetic instructions yielding robust models, and, most importantly, that using as backbone an instruction-tuned model outperforms using a base non-instructed model, and improved results when scaling up. Using Llama 3.1 instruct 70B as backbone our model comes near frontier models of much larger sizes for Basque, without using any Basque data apart from the 1.2B word corpora. We release111Repository: https://github.com/hitz-zentroa/latxa-instruct code, models, instruction datasets, and human preferences to support full reproducibility in future research on low-resource language adaptation.
booktabs
Instructing Large Language Models for Low-Resource Languages:
A Systematic Study for Basque
Oscar Sainz Naiara Perez Julen Etxaniz Joseba Fernandez de Landa Itziar Aldabe Iker García-Ferrero Aimar Zabala Ekhi Azurmendi German Rigau Eneko Agirre Mikel Artetxe Aitor Soroa HiTZ Center - Ixa, University of the Basque Country UPV/EHU {oscar.sainz,a.soroa}@ehu.eus
1 Introduction
Large Language Models (LLMs), particularly open models, remain predominantly English-centric, with limited coverage for the vast majority of the world’s languages. Despite recent efforts to incorporate additional languages during the pretraining of open LLMs, significant performance disparities still persist. Even the latest instruction-tuned models demonstrate markedly degraded capabilities when handling low-resource languages (Grandury et al., 2025). Critically, the English-focused nature of post-training processes has widened the performance gap between languages when comparing base and instruction-tuned models.
To overcome these limitations, open models can be adapted to new languages through continued training with limited resources Etxaniz et al. (2024b). In the case of instruction-tuned models in particular, various efforts have emerged that typically follow a sequential approach Ouyang et al. (2022): first adapting the base model through continued pretraining, then performing instruction tuning. While this multi-step process has become standard practice, little exploration has investigated alternative adaptation strategies. We question whether instruction-following capabilities could be directly transferred to new languages without dedicated instruction data, and whether instruction-tuned models could be adapted through continued pretraining, similar to base models.
Specifically, this work systematically explores diverse strategies beyond the conventional pipeline for developing instruction-tuned models for low-resource languages, seeking to identify optimal adaptation paths for Basque as our primary case study (see Fig. 1). We deliberately constrain our exploration to resources either readily available or creatable using open models, avoiding reliance on distillation from commercial state-of-the-art systems. While our investigation focuses on a single language, our findings likely generalize to many similarly-resourced languages: Basque represents an ideal test case, ranking approximately 50th in Common Crawl with a presence approximately times smaller than English,222https://commoncrawl.github.io/cc-crawl-statistics/plots/languages.html and notably lacking pre-existing instruction datasets. This resource profile mirrors challenges faced by numerous other low-resource languages worldwide.
In addressing this research question, we further confront a key challenge in instruction-following LLM assessment: automated metrics often miss capabilities that matter to users Chiang et al. (2024). Thus, we developed an evaluation framework combining traditional benchmarks with a crowdsourced LLM arena, where we mobilized the Basque-speaking community in a large-scale evaluation effort that gathered over preference annotations from participants. This initiative constitutes the largest human evaluation effort for a low-resource language to date.
Through this evaluation, our systematic exploration produced three key insights for developing instruction-tuned models in low-resource languages: (1) target language corpora is essential for performance—models lacking exposure to plain Basque text showed degradation regardless of other techniques; (2) while both monolingual and bilingual instruction datasets showed benefits, the latter produced consistent results across benchmark and human evaluations; and (3) starting from an instruction-tuned English model outperformed the approach of a base model learning to follow instructions, challenging the standard pipeline applied to low-resource languages.
In addition to these primary findings, our work makes the following contributions to the field: (4) the first release of an instruction-tuned family of LLMs specifically for Basque, in sizes of 8B and 70B parameters, the latter of which proved competitive with GPT-4o and Claude Sonnet in the arena; (5) the release of large-scale, synthetic instruction-tuning datasets in English and Basque; and (6) the release of the first preference dataset in Basque, containing real user prompts, model responses, and preference annotations that could support future preference alignment research. Through these contributions, we aim to advance the state of language technology for Basque while establishing methodologies applicable to other low-resource languages.
2 Related Work
Research on developing LLMs for under-resourced languages has explored various approaches, with varying degrees of success. Initial attempts to develop models from scratch for specific low-resource languages have proven challenging due to limited training data. Multilingual model development has emerged as a more promising strategy, with researchers leveraging cross-lingual transfer learning to improve performance Scao et al. (2023); Le Scao et al. (2022); Shliazhko et al. (2024). The most effective approach to date involves continued pre-training of existing multilingual models, which allows for language-specific adaptation while benefiting from the rich linguistic representations of larger training corpora Etxaniz et al. (2024b); Luukkonen et al. (2023); Tran et al. (2024). While progress has been made in developing these base models, optimal methods for instructing and fine-tuning them for under-resourced languages remain largely unexplored Gonzalez-Agirre et al. (2025); Martins et al. (2025); Üstün et al. (2024).
Instruction-tuning for under-resourced languages has explored various approaches to overcome the scarcity of native instruction data. Different studies leverage either English-centric pretrained models or multilingual models as pivot architectures for cross-lingual transfer Purason et al. (2025). Regarding the data, researchers have explored incorporating multilingual instruction datasets that include limited coverage of lower-resourced languages Shaham et al. (2024); translating existing English instruction sets into target languages either automatically or with human verification Joshi et al. (2025); Zosa et al. (2025); and applying data augmentation techniques like back-translation, language-specific prompting, and template-based instruction generation to expand limited resources Li et al. (2024). Additionally, cross-lingual in-context learning has shown interesting results Cahyawijaya et al. (2024).
Regarding Basque language adaptation, two significant studies have been conducted. Etxaniz et al. (2024b) developed Latxa by adapting Llama 2 models through continued pretraining. Meanwhile, Corral et al. (2025) created Llama-eus by adapting Llama 3.1 and subsequently performing both instruction tuning and preference alignment using machine-translated data, adhering to widely accepted methodologies. However, the former focuses solely on foundation models, without considering instruction-tuned models, while the latter implements only a single strategy for instructing an adapted language model. In contrast, this work explores multiple strategies and combinations systematically to effectively instruct (or adapt) a language model for Basque.
3 Resources
Instructing LLMs typically relies on two components: base (or foundational) LLMs and instruction datasets. For non-hegemonic languages, obtaining instruction datasets can be very challenging, particularly in low-resource language scenarios. In the case of Basque in particular, there are no manually generated, or even good quality automatically generated, large sets of instruction-answer pairs. Consequently, as shown in Fig. 1, our available resources are constrained to corpora on the target language, and base and instruct models for high-resource languages. From these limited resources, we derive the necessary components to create Basque instruction-tuned models through strategic combinations of synthetic data generation and model adaptation. In the following sections, we describe these seed resources and derivations.
3.1 Basque corpora
For the pre-training data, we have leveraged the corpora used to train Latxa, the first family of LLMs trained specifically for Basque Etxaniz et al. (2024b). This corpus comprises 4.3M of high-quality documents in Basque, roughly 3.5B Llama 3.1 tokens. Among the sources, it contains high-quality news data extracted using ad-hoc scrapers Artetxe et al. (2022), Wikipedia333The 20231101 dump corresponding to November 2023. and sources based on Common Crawl such as CulturaX Nguyen et al. (2024), Colossal OSCAR Abadji et al. (2022) and HLPT v1.1 de Gibert et al. (2024). This corpus comes normalized, deduplicated and filtered. The data is publicly available in the HuggingFace hub.444hf.co/datasets/HiTZ/latxa-corpus-v1.1 We will henceforth refer to this corpus as Corpus EU.
3.2 Backbone models
As our base LLM (i.e., models that have not been fine-tuned to follow chat-style instructions) we use Llama 3.1 Grattafiori et al. (2024). Llama 3.1 is a publicly available model widely adopted by the community due to its strong performance across English and other high-resource languages. We refer to this model as BaseEN throughout the paper. In addition, following Etxaniz et al. (2024b), we train a new Latxa model based on Llama 3.1, which we denote as BaseEU. For the instruction-tuned models, we adopt a similar strategy and use the instruction-following version of Llama 3.1, which we refer to as InstructEN.
3.3 Instruction Sampling and Translation
Existing (English) instruction datasets rely on either high-quality, manually crafted instructions and responses (e.g., No Robots),555hf.co/datasets/HuggingFaceH4/no_robots fully automatically generated instructions and responses Ding et al. (2023); Ge et al. (2025), or a combination of both, such as manually written prompts paired with automatically generated responses Zhao et al. (2024). Using any of these datasets would introduce an additional confounding factor into our analysis (namely, knowledge distilled from a powerful LLM), which could lead a model trained on such data to outperform our InstructEN, thus introducing noise into our evaluation. This would raise a separate research question that falls outside the scope of this paper: what is the best (combination of) instruction dataset(s) to train a model on? In the case of Basque, however, there is no publicly available set of instructions. The following paragraphs detail the process of generating the instructions for each language.
English instructions.
To avoid external influences, we instead sample instructions directly from our InstructEN model. We generate the English instructions following Xu et al. (2025). Using this technique, we conditioned InstructEN to generate instructions of different types and tasks: general-purpose, code, math, arithmetic and translation. We generated a total of 4M English instructions, however, after a hyperparameter search, we found out that using just 1M instructions yielded better results overall. We share more details and examples of the process in Appendix A.
Basque instructions.
We translated instructions sampled from InstructEN using few-shot prompting with Base EU. Existing machine translation systems for the English–Basque language pair (e.g., NLLB Costa-jussà et al. (2022)) are primarily trained on sentence-level textual data and often struggle with more complex inputs, including selectively translating natural language content embedded within code snippets. By leveraging an LLM like Base EU, which has been exposed to diverse data types, we obtained higher-quality translations for this setting. Moreover, using a model trained within our own experimental framework allows us to avoid introducing external factors into our pipeline. More details about the process and prompts used to translate the instructions are given in Appendix A.
4 Experimental Setup
We formalize our experimental setup as follows. Let be the set of backbone models and be the set of binary variables indicating whether to use Basque corpora, English instructions, and/or Basque instructions. The space of possible configurations is thus , where is the power set of , yielding theoretical combinations. Note that we explore training strategies that leverage both raw text and instruction data simultaneously. From the total of 24 combinations, we exclude redundant configurations where a model is retrained on data it was originally trained with. The resulting set of distinct instruction-tuned model variants therefore comprises 18 configurations: the original Llama 3.1 Instruct 8B (i.e., Instruct EN) and 17 new 8B-sized models. Table 3 in Appendix B provides a complete account of all model variants and their shorthand names. Additionally, we trained a 70B model following the configuration that performed best in preliminary benchmark evaluations.
Regarding the baselines, the primary baseline in our analysis is the Instruct EN model, as it is the only backbone capable of following instructions. However, since we examine the effect of each variable in individually, the specific points of comparison used vary across cases. For additional context, we also evaluate two proprietary models known for their strong performance in Basque:666We did not include the instructed Llama-eus Corral et al. (2025) in our evaluation, as it was not publicly available at the time of experimentation. OpenAI’s GPT-4o777gpt-4o-2024-11-20 and Anthropic’s Claude 3.5 Sonnet.888claude-3-5-sonnet-20241022
5 Evaluation
We employed two complementary evaluation approaches to assess the impact of each instruction-tuning recipe. On the one hand, we used a selection of static benchmarks that evaluate specific model capabilities and knowledge through standardized tests. On the other hand, we conducted human evaluations through A/B testing (arena style) to capture qualitative aspects of model performance.
5.1 Static Benchmarks
We selected benchmarks that are close to real use cases, from a varied range of categories: reading comprehension, common sense, linguistic proficiency, knowledge, maths & reasoning, and bias. For each benchmark, where possible, we evaluated the Basque, English, and Spanish versions to facilitate the analysis of language-specific tradeoffs for each fine-tuned model variant. We included Spanish due to its close sociolinguistic relationship with Basque and to assess how our instruction-tuning methods affect a third language not directly targeted by our experiments. Thus, we evaluated a total of 29 benchmarks, as detailed in Section C.1.
For conducting these evaluations, we relied on Eleuther AI’s LM Evaluation Harness Biderman et al. (2024). Most datasets are framed as multiple-choice problems where models’ answers are determined by selecting the option with the highest log probability. For generative tasks, answers are directly sampled from the model. To provide models with contextual examples, our evaluations employed a few-shot setting. All results are measured for accuracy following standard, public implementations.999We implemented some benchmarks that are not currently available, and which will be contributed to the main repository (github.com/EleutherAI/lm-evaluation-harness) upon publication. Refer to Section C.1 for details.
When evaluating proprietary models, we cannot directly compute log probabilities because we have no access to model weights. This limitation restricts our evaluation to only those benchmarks implemented as explicit letter-choice questions (A, B, C, …) and the free-form generative task MGSM, excluding benchmarks that require comparing verbalized option likelihoods. For the compatible multiple-choice benchmarks, we prompt models to output a single letter as their answer, using the same prompts and few-shot examples as with open models to maintain comparability.
5.2 Human Evaluation: Arena
Unlike static benchmarks, which rely on fixed datasets and automatic metrics, arena-style evaluations are better suited for assessing open-ended text generation, where subjective quality judgments play a central role. In this section, we first describe our implementation of the arena framework, including details on participants and evaluation conditions. For additional details about the human evaluation, refer to Section C.2, where we describe our infrastructure and introduce the Bradley–Terry model Bradley and Terry (1952), which we use to infer a model ranking from the collected pairwise preferences.
To gather human preferences for our evaluation, we organized a community-driven initiative. This collaborative effort ran for 14 days and attracted approximately 1,680 participants, resulting in a total of 12,890 preference annotations. The event was open to any Basque speaker, regardless of their proficiency level. Participants were required to register their educational background and language proficiency before contributing. Once registered, users could submit prompts and compare model responses. Fig. 2 shows that the majority of the participants—and, therefore, the preferences—have a bachelor or superior education and a high or native language proficiency level.
In the annotation process, participants evaluated pairs of model responses by making a three-way choice (i.e., prefer model ‘A’, prefer model ‘B’, or consider them tied) across two dimensions: content quality and linguistic quality. Linguistic quality was considered as a separate dimension because not all models produce fluent and sound Basque—an uncommon issue among high-resource languages. In cases where participants’ judgments were contradictory between the dimensions, a third question about overall quality was presented to determine the final choice.
It merits mention that we offered prizes based on user activity to encourage widespread participation and maximize the number and diversity of collected preferences. While this strategy succeeded in increasing engagement, it also attracted malicious or dishonest users who prioritized quantity over quality to win rewards. We proactively identified and banned such actors using a combination of heuristics and manual review.
6 Results
In the following paragraphs, we disclose the effect of each component: backbone models, Basque corpora and bilingual instructions. For both the benchmark and the human evaluation, main results are presented in Fig. 3. Complete results for all tested models, broken down by benchmark and language, along with detailed arena evaluation results, can be consulted in Appendix D.
6.1 The Impact of Basque Corpora
We begin by analyzing the influence of Corpus EU, as it is intuitively the most critical resource for teaching a new language to an LLM. Fig. 3 confirms that this intuition aligns well with empirical results. In both human evaluations (Fig. 3(a)) and benchmark scores (Fig. 3(b)), models trained on Corpus EU achieved significantly better performance. The advantage is especially pronounced—up to 12 points in accuracy and over 300 points in arena score—when no other Basque signal (e.g., Basque instructions) is included. However, for models that already use Base EU as their backbone, additional exposure to Corpus EU offered little benefit and was sometimes even detrimental.
We conclude that using target language corpora is highly beneficial and possibly essential for training an instruction-tuned LLM in our low-resource language. Therefore, the following analyses will focus exclusively on the variants trained with Corpus EU.
[…]
| BertaQA Global | 67.10 | 74.62 | 73.54 | 72.99 | 83.53 | 87.42 | 93.52 | 91.01 |
| BertaQA Local | 44.97 | 65.23 | 66.07 | 65.57 | 53.51 | 77.71 | 80.45 | 74.83 |
| EusProficiency | 34.13 | 52.83 | 52.06 | 52.35 | 43.59 | 68.00 | 81.60 | 74.25 |
| EusReading | 49.72 | 59.66 | 62.78 | 61.93 | 72.16 | 78.98 | 87.39 | 84.38 |
| EusTrivia | 45.01 | 61.05 | 62.33 | 62.10 | 62.51 | 74.17 | 84.60 | 80.70 |
| EusExams | 46.21 | 56.00 | 56.01 | 56.23 | 63.28 | 71.56 | 82.68 | 79.17 |
| MGSM | 45.60 | 54.00 | 46.40 | 50.80 | 76.40 | 80.00 | 85.20 | 79.20 |
| MMLU | 50.37 | 57.04 | 52.96 | 56.30 | 68.52 | 68.89 | 79.63 | 76.66 |
| Benchmark Avg | 50.78 | 62.27 | 61.51 | 62.36 | 68.07 | 77.53 | 85.48 | 81.45 |
[…]
| Arena Language | 783 (-12,+12) | 1036 (-10,+11) | 1034 (-10, +8) | 1038 ( -8,+10) | - | 1083 (-13,+13) | 1082 (-11,+11) | 1093 (-10,+12) |
| Arena Global | 722 (-17,+19) | 1038 (-13,+13) | 1050 (-11,+14) | 1050 (-14,+13) | - | 1141 (-11,+15) | 1153 (-21,+13) | 1188 (-17,+13) |
[…]
6.2 The Impact of Instruction Data
We analyze the effect of instruction data by comparing variants trained with no instructions, English-only instructions, Basque-only instructions, and a combination of both.
Starting with the question of whether to use instructions at all, we focus on Instruct EN-based variants. Human evaluation results (Fig. 3(a)) clearly show that incorporating instructions, regardless of language, helps mitigate catastrophic forgetting and improves arena scores by nearly 100 points. In contrast, benchmark results (Fig. 3(b)) show only marginal gains, particularly when using Basque instructions. This discrepancy highlights the limitations of static benchmarks and underscores the value of human or text generation-based evaluation.
When comparing instruction languages, we observe a general trend: English instructions tend to yield better results across both evaluation methods. However, there are exceptions. For example, models based on Instruct EN perform comparably or slightly better with Basque instructions in human evaluations, while Base EU-based models perform similarly on benchmarks. Notably, combining English and Basque instructions consistently produces the best results across most scenarios.
Although the improvements are not always significantly better, we conclude that including instructions in both languages results in more robust models, achieving stronger performance regardless of the backbone. Consequently, the remainder of our analysis will focus on models trained with bilingual instruction data.
6.3 The Impact of Backbone Models and Curriculum Learning
By analyzing models trained from different backbones, we explore various curriculum learning strategies: (i) teaching the language first and then instruction-following, (ii) teaching instruction-following in English first and then the target language, or (iii) learning everything simultaneously.
Language first vs. simultaneously.
When comparing models based on Base EN (learning everything together) with those based on Base EU (learning the language first), we observe no significant difference in performance. Interestingly, the Base EU variant without access to Corpus EU achieves performance nearly identical to the Base EN variant with access to Corpus EU, across all instruction settings. This suggests that teaching the language in a separate pre-training step offers no measurable advantage. From this, we conclude that there is no compelling reason to separate language acquisition from instruction tuning.
Instructions first vs. simultaneously.
While instruction tuning pipelines are often complex and multi-staged Lambert et al. (2025), our approach adopts a simpler structure. Previous work by Xu et al. (2025) showed that models initialized from Base EN and trained with sampled instructions from Instruct EN perform comparably to Instruct EN. However, our findings indicate that this strategy does not transfer well to low-resource languages. As shown in Fig. 3, models based on Instruct EN consistently outperform those based on Base EN, both in human evaluations and benchmarks. These results support the conclusion that starting from a well-instructed English backbone yields better performance than learning everything from scratch.
6.4 The Impact of the Scale
Based on our previous analyses, we scaled up the variant that performed best in preliminary benchmark evaluations.101010We did this before running the arena as we wanted to include a 70B model in the human evaluation. Table 1 shows the results for some multiple-choice benchmarks and arena scores. On the one hand, we have the three best-performing 8B variants and the baseline. On the other hand, we present the results for the 70B best-performing variant and the baseline. Finally, we also show the performance of Claude 3.5 Sonnet and GPT-4o.
Scaling to bigger sizes.
We analyzed the effect of our language adaptation process when training a larger model. Despite the results of the 70B Instruct EN baseline being significantly better than the 8B counterpart (even surpassing the Instruct EU variants), we observe that our language adaptation step still had similar improvements to those obtained with 8B sized models—almost 10 accuracy points gain on average. The biggest gains are observed in local knowledge and language proficiency, the only benchmarks where the 70B Instruct EN underperforms the best 8B variant.
Comparing to the State of the Art.
When compared to the leading commercial models in Basque, our best model falls slightly behind in most benchmarks except for BertaQA Local and MGSM, where our model performs better than GPT-4o—particularly in local knowledge about the Basque Country. Regarding the arena score, our best model is on par with the commercial models in perceived linguistic quality, but falls behind the best model in the content quality and global scores. Interestingly, Claude 3.5 Sonnet outperforms GPT-4o in all benchmarks, but the latter gets a higher score in the arena. Our best model being worse than SotA despite focusing on Basque might be related to the weaker backbone model we used, as Instruct EN is overall a weaker model. Using a stronger and larger backbone model in the future could lead to results that match the SotA models in benchmarks and arena score.
7 Analysis and Discussion
In this section, we provide additional analysis and discussion of our results. First, we focus on the correlation of our two main evaluation methods. Then, we measure the trade-off between Basque and other languages. Finally, we analyze the safety and biases of our models.
Benchmark–Arena correlation.
We find a strong correlation between Basque benchmark performance and human preferences in the arena, with Spearman’s rank correlation coefficients exceeding 0.80 across all arena dimensions (see Table 2), which suggests that automated Basque benchmarks may provide a reliable proxy for human evaluation in future research. English benchmarks display positive but non-significant correlations, while Spanish benchmarks show virtually no correlation with our Basque arena results. It must be noted that these correlations are not directly comparable across languages due to differences in benchmark composition. Detailed correlations by specific dataset can be found in Section D.1.
width=colspec=lX[2,r]X[3,l]X[2,r]X[3,l]X[2,r]X[3,l],
cells=font=,
row1=font=,
column1=font=,
column1=rightsep=3pt,
column2,4,6=rightsep=2pt,
column3,5,7=leftsep=1pt, rightsep=3pt,
rowsep=0pt
[…]
Bench EN 0.25 P = 0.32 0.23 P = 0.37 0.27 P = 0.29
Bench ES -0.02 P = 0.95 -0.09 P = 0.72 -0.04 P = 0.89
[…]
Trade-off between Basque and other languages.
Fig. 4 shows performance changes across languages relative to each backbone model on the multilingual benchmarks Belebele, MGSM, MMLU, and XStoryCloze. We observe that Instruct EN-based models exhibit a clear trade-off: improvements in Basque come at a cost of decreased performance in English and Spanish, suggesting a competitive relationship between languages in the model’s parameter space. In contrast, greater flexibility for multilingual adaptation is observed in Base EN models, which improve across all three languages (though from a lower absolute performance baseline). Base EU models show moderate changes with gains primarily in Spanish and English rather than Basque. As observed previously, Corpus EU consistently yields the largest performance gains for the target language. Among Instruct EN variants, the configuration with Corpus EU and Instructions EN achieves the most Basque improvement and the least regression in other languages. Despite these adaptation strategies, models still perform better in English and Spanish on equivalent tasks, with the exception of culturally-specific knowledge as evidenced by the results on the BertaQA dataset (see complete results in Section D.2).
Safety and bias.
This paper does not address the alignment step typically required to reduce biases and prevent unsafe responses in production-ready LLMs. However, the backbone model Instruct EN is already aligned, and we expect variants based on this backbone to retain safety-related behaviors. To verify this, we conducted a small-scale experiment to assess the safety and bias characteristics of our models after the language adaptation (see Appendix E). Results indicate that the models preserve most of the safety and bias alignment properties of Instruct EN in English, and that these properties effectively transfer to Basque, with no observed degradation after adaptation.
8 Conclusions
This systematic study on instruction-tuning LLMs for Basque reveals several key strategies for low-resource language adaptation. We found that target language corpora are essential for effective learning. Employing bilingual (English and Basque) synthetic instructions yielded the most robust models whereas English-only instructions remain competitive. Crucially, starting from an instruction-tuned English model and adapting it to Basque proved more effective than training a base model for both language and instruction-following, or pre-training for language separately before instruction tuning. Our work contributes new Basque models, open instruction and human preference datasets, and methodological insights to guide future low-resource LLM development.
In the future, we plan to extend the exploration using instructions created by humans. The preference data we release can also be used to align the models. Scaling to stronger backbones will also lead to better results that could match the SotA commercial models in Basque. The strong correlation observed between aggregated Basque benchmarks and human evaluations also suggests a path for more efficient proxy evaluations.
Limitations
This study presents a systematic analysis aimed at identifying the most effective method for developing an instruction-tuned model for a low-resource language. However, due to the combinatorial nature of such analyses, we had to constrain certain dimensions of our exploration, as adding any additional axis would effectively double the amount of work required, including costly human evaluations.
Our first limitation is the choice of language. We focused on Basque, primarily because it is a low-resource language with just enough available data to enable language adaptation of base models Etxaniz et al. (2024b). While some other languages have significantly fewer resources than Basque, our conclusions may not fully generalize to those more extreme low-resource scenarios.
The second limitation is the choice of model family. We conducted all experiments using the Llama 3.1 family as the backbone. This decision was motivated by its widespread adoption and its existing ability to generate text in Basque, although often with substantial linguistic errors. Evaluating more recent or higher-performing models could slightly influence our findings. However, to the best of our knowledge, there is currently no open model family capable of producing linguistically correct Basque.
Finally, this study primarily focuses on the initial instruction-tuning phase. While we did collect preference data, we did not extend our analysis to include preference alignment techniques. Including this additional phase would again have doubled the experimental workload and human evaluation requirements.
Acknowledgements
This work has been partially supported by the Basque Government (Research group funding IT1570-22 and IKER-GAITU project), the Spanish Ministry for Digital Transformation and of Civil Service, and the EU-funded NextGenerationEU Recovery, Transformation and Resilience Plan (ILENIA project, 2022/TL22/00215335). The models were trained on the Leonardo supercomputer at CINECA under the EuroHPC Joint Undertaking, project EHPC-EXT-2024E01-042. Julen Etxaniz and Ekhi Azurmendi hold a PhD grant from the Basque Government (PRE_2024_2_0028 and PRE_2024_1_0035, respectively).
Appendix A Synthetic Instructions generation
A.1 English instructions
To generate the English synthetic instructions, we followed a technique called Magpie Xu et al. (2025). Briefly, it consists of letting the model generate text starting from the user’s prompt instead of the assistant’s response. See, for instance Fig. 5(a), where the model is asked to continue with the chat template immediately after the user header. We defined 5 different prompts to generate different kinds of instructions (see Fig. 5). We generated instructions sampling from the model using different 10 temperature values, ranging from 0.8 to 1.2. After generating the instructions, we apply the following filters:
1.
Remove duplicates: we remove those instructions that are exactly the same.
2.
Repetitive prompts or responses: we remove those instances that contain words that repeat more than 100 times.
3.
Poor quality instructions or responses: we remove those instances that were classified as poor quality by the model itself.
4.
Unfinished instructions: we noticed that some instructions ended with ":", meaning that the instructions was incomplete. We removed those as well.
As we experimented with models of two sizes and did not want to have any external influence in our experiments, we generated the instructions twice, once per model size. We later used the instructions generated from the 8B model to train the 8B models, and those from 70B to train the 70B model. The total size in tokens of the English instructions used for training is 350M tokens. We are releasing111111To be released upon acceptance. the generated instructions (filtered and unfiltered) for replicability purposes.
A.2 Basque (translated) instructions
As discussed in Section 3.3, we generated our Basque instructions by translating the English instructions using Base EU. The prompt we used to perform the translation is shown in Fig. 6. As for the few-shot examples, we included 2 conversations of general instructions and 1 conversation for math, arithmetic, code and translation instruction types. These examples were sampled from the English instructions and post-edited manually. The exact examples are released in the GitHub repository.
Using a base LLM introduces several challenges when generating structured data. To mitigate issues with malformed outputs, we constrained the LLM’s generation to adhere to a conversational JSON schema. Additionally, a non-negligible portion of the outputs (approximately 1%) exhibited repetition, either in the question or the response. To prevent these repetitive examples from being incorporated into the instruction-tuned model, we applied simple heuristics based on span frequency to identify and filter them out. The total size in tokens of the Basque instructions used for training is 503M tokens. We are releasing the translated instructions for replicability purposes.
A.3 Examples
Appendix B Training details
colspec=X[3]X[2]X[2]X[2],
cells=l,font=,
row1=c,font=,
row2=l,font=,
column1=l,font=,
colsep=2pt,
rowsep=0pt
[…]
Base EN Base EU Instruct EN
Instructions EN
[…]
BAS EU I EN
[…]
Instructions EU
BAS EN I EU
BAS EU I EU
INS EN I EU
Instructions EN+EU
BAS EN I EN+EU
BAS EU I EN+EU
INS EN I EN+EU
Corpus EU
[…]
INS EN C EU
Corpus EU + Instructions EN
BAS EN C EU I EN
BAS EU C EU I EN
INS EN C EU I EN
Corpus EU + Instructions EU
BAS EN C EU I EU
BAS EU C EU I EU
INS EN C EU I EU
Corpus EU + Instructions EN+EU
BAS EN C EU I EN+EU
BAS EU C EU I EN+EU
INS EN C EU I EN+EU
[…]
colspec=lrrrr,
cells=font=,
row1=c,font=,
rowsep=0pt
[…]
Any backbone + I EU 64Gb A100 128 264.5h 94.26Kg
Any backbone + I EN+EU 64Gb A100 128 456.5h 162.69Kg
Any backbone + C EU 64Gb A100 128 1,730.1h 205.54Kg
Any backbone + C EU I EN 64Gb A100 128 1,932.8h 688.86Kg
Any backbone + C EU I EU 64Gb A100 128 2,016.0h 718.50Kg
Any backbone + C EU I EN+EU 64Gb A100 128 2,327.5h 829.53Kg
[…]
Hardware and carbon footprint.
Table 4 summarizes the training costs of our experiments (see Table 3 for the variants nomenclature). In total, we trained 19 models: one Base EU, seventeen Instruct EU variants, and one 70B Instruct EU. Due to unforeseen circumstances, the Base EU and Instruct EU models were trained using different frameworks and infrastructure. The Base EU model was trained with NeMo Harper et al. (2024) on 64GB H100 GPUs provided by MareNostrum 5. In contrast, all Instruct EU variants were trained using Fully Sharded Data-Parallel Zhao et al. (2023) on 64GB A100 GPUs provided by CINECA Leonardo. The total compute time across all experiments amounted to 39,879.1 GPU hours, which corresponds to 4,729.11 kg CO2eq, based on carbon intensity estimates from ElectricityMaps.121212At the time of the experiments: 0.297 kg/kWh for Italy and 0.157 kg/kWh otherwise, according to https://www.electricitymaps.com/.
colspec = lcc,
cells=font=,
row1 = font=,
rowsep=0pt
[…]
Total Batch Tokens 2M 4M
[…]
FSDP Sharding Strategy HYBRID FULL
[…]
Hyperparameters.
Table 5 outlines the key hyperparameters used during training. Both 8B and 70B model variants were trained with consistent configurations in terms of sequence length, batch size, optimizer settings, and learning rate schedules. The main differences lie in the number of GPUs used and the FSDP sharding strategy, which was adjusted to better accommodate the increased memory and compute demands of the larger model. These hyperparameters were optimized using 8B model variant of Base EU, and the initial iteration on the instructed variants. We found this configuration to robustly perform across all the configurations.
Appendix C Evaluation details
C.1 Static Benchmarks
Our evaluation framework comprises a total of 29 benchmarks across three languages: 15 in Basque, 10 in English, and 4 in Spanish. These benchmarks span six categories designed to test different aspects of model capabilities:
Common sense: XStoryCloze Lin et al. (2022), a multilingual version of the original StoryCloze Mostafazadeh et al. (2017) dataset testing narrative understanding; and PIQA Bisk et al. (2020), which assesses physical common sense through everyday tasks. PIQA’s translation to Basque is available through IberoBench Baucells et al. (2025).
Miscellaneous knowledge: BertaQA Etxaniz et al. (2024a), which tests knowledge of local Basque culture versus global topics; EusTrivia and EusExams from the Latxa suite Etxaniz et al. (2024b); and a subset of MMLU Hendrycks et al. (2021), manually translated to Basque Corral et al. (2025) and Spanish.131313https://huggingface.co/datasets/openai/MMMLU
Except MGSM, the datasets are framed as multiple-choice problems where models’ answers are determined by selecting the option with the highest log probability. MGSM is implemented as a generative task where an answer is directly sampled from the evaluated model and matched against a reference answer. We specifically chose the native chain-of-thought scenario. To provide models with contextual examples, our evaluations employed a 5-shot setting, except for BBQ and BasqBBQ which use 4 examples.
C.2 Arena Details
Guidelines for arena participants.
Fig. 10 contains the information and instructions that were given to the human annotators who participated in the community-driven arena initiative. All the data collected through this initiative was properly anonymized prior to publication. Note that the actual information was provided in Basque, while here we show a translation to English.
Arena infrastructure.
On the infrastructure side, we used vLLM Kwon et al. (2023) to serve all model pairs and the baseline. We deployed a total of 18 endpoints for 8B models and one endpoint for a 70B model, running on nine and two A100 80GB GPUs, respectively. For the frontend, we developed a lightweight Gradio Abid et al. (2019) interface that allowed users to enter prompts, view model responses, and indicate their preferences based on content quality, language quality, or, in cases where no clear winner emerged, overall quality. To ensure a fair comparison across models, all models were given the same system prompt (shown in Fig. 9) and the same hyperparameters: 0.9 temperature and 0.95 top-p.
Bradley–Terry model.
The Bradley–Terry model Bradley and Terry (1952) provides a principled probabilistic framework for aggregating pairwise preferences into a global ranking over models. Let denote the set of models under evaluation. The model assigns a latent preference strength to each model . The probability that model is preferred over in a pairwise comparison is given by:
Given a dataset of observed pairwise outcomes the parameters are estimated using Maximum Likelihood Estimation (MLE). To facilitate interpretation, we apply zero-mean centering, treating the scores as deviations from the average model. The final scores for each model are then computed as:
By using a scaling factor of 400, we ensure that the scores are interpretable in a manner consistent with the online ELO rating system. Thanks to its score stability and the assumption that model performance remains constant over time, the Bradley–Terry scoring system has become a widely adopted method for ranking LLMs—particularly since its introduction in the Chatbot Arena.141414https://lmarena.ai
Appendix D Detailed Results
D.1 Benchmark–Arena Correlation
Fig. 11 presents a detailed analysis of Spearman’s rank correlation between individual benchmark performance and arena scores across different benchmark languages (Basque, English, Spanish) and arena preference dimensions (global, content, language).
As discussed in Section 7, Basque benchmarks show strong and consistent correlations with human preferences across all arena dimensions. This correlation is particularly pronounced for specific benchmarks including EusProficiency, EusTrivia, and EusExams, which are interestingly the datasets that were natively constructed in Basque, rather than translated from existing English benchmarks. BertaQA Local also shows notable correlations in English, likely reflecting the types of culturally-specific questions that users posed in the arena evaluation.
Notably, bias benchmarks (Basq)BBQ show virtually no correlation with human preferences, regardless of the language. This disconnect may stem from the absence of specific fairness assessment guidelines in this community-driven evaluation. Participants likely neither probed for the types of biases BBQ is designed to measure nor prioritize these aspects in their preference judgments.
D.2 Benchmark Results
colspec=l*15X[c],
cells=font=,
row1=c,font=,
column1=font=,
colsep=3pt,
rowsep=0pt
[…]
+ I EU 36.35 61.95 69.44 67.39 42.98 84.06 48.39 35.42 44.89 45.42 30.80 50.37 61.44 62.14 52.93
+ I EN+EU 38.28 63.50 72.00 68.76 42.77 84.25 48.41 36.19 45.17 46.27 37.20 52.04 62.09 62.84 54.27
+ C EU I EN 39.42 64.81 73.00 70.99 63.49 91.28 51.92 48.66 56.82 56.33 47.20 51.85 67.32 69.36 60.89
+ C EU I EU 38.14 67.00 71.78 71.24 62.65 92.39 49.88 47.20 46.02 57.78 39.20 50.00 65.63 66.71 58.97
+ C EU I EN+EU 40.10 68.60 74.89 72.74 63.79 92.50 52.78 47.65 53.12 59.59 48.40 51.48 64.87 66.84 61.24
[…]
+ I EN 40.36 63.93 72.22 70.94 63.20 90.11 52.21 48.31 59.09 57.78 45.60 52.22 65.85 70.55 60.88
+ I EU 39.25 67.09 73.56 71.78 61.59 92.11 52.38 47.78 49.43 57.14 35.60 53.70 64.92 67.31 59.55
+ I EN+EU 38.31 66.92 74.56 72.58 62.23 91.22 53.04 48.96 53.98 58.95 44.80 56.30 64.27 67.70 60.99
+ C EU I EN 40.96 66.12 61.22 71.49 63.83 91.72 50.67 47.19 47.16 57.49 43.60 52.59 66.83 70.35 59.37
+ C EU I EU 37.97 67.42 71.33 71.24 62.52 91.83 51.76 46.31 52.84 57.32 37.60 52.22 65.90 67.31 59.54
+ C EU I EN+EU 39.16 68.43 71.67 71.70 64.13 92.06 52.61 48.09 53.69 58.43 40.80 54.44 66.23 68.63 60.72
[…]
+ I EU 38.65 63.85 78.00 69.57 42.98 83.67 51.80 38.36 50.57 45.20 27.20 55.56 62.64 64.73 56.76
+ I EN+EU 39.59 64.65 79.22 70.40 43.10 84.00 51.43 38.69 52.56 52.80 35.20 54.81 62.85 64.13 57.51
+ C EU 37.97 65.70 77.33 73.87 66.33 92.67 55.05 52.12 58.24 61.40 48.40 51.85 66.99 70.28 62.73
+ C EU I EN 41.38 66.79 80.00 74.62 65.23 91.39 56.00 52.83 59.66 61.05 54.00 57.04 67.32 71.34 64.19
+ C EU I EU 40.44 69.15 81.44 73.54 66.07 91.83 56.01 52.06 62.78 62.33 46.40 52.96 66.01 71.01 63.72
+ C EU I EN+EU 39.85 70.16 83.00 72.99 65.57 92.28 56.23 52.35 61.93 62.10 50.80 56.30 65.69 69.56 64.20
[…]
+ C EU I EN 55.12 77.57 91.00 87.42 77.71 92.11 71.56 68.00 78.98 74.17 80.00 68.89 70.75 77.83 76.51
[…]
GPT-4o - - 92.89 91.01 74.83 - - 74.25 84.38 80.70 79.20 76.67 - - -
[…]
colspec=l*15X[c],
cells=font=,
row1,2=c,font=,
column1=font=,
colsep=3pt,
rowsep=0pt
[…]
ARC C ARC E Bele BQA G BQA L MGSM MMLU PIQA XSC Avg Bele MGSM MMLU XSC Avg
[…]
+ I EU 54.95 83.63 86.78 73.95 47.38 67.60 63.70 80.63 82.46 70.73 80.67 58.40 57.78 74.12 67.74
+ I EN+EU 54.82 81.42 87.61 74.46 47.91 78.00 61.30 80.25 82.40 72.52 82.00 67.60 58.89 75.28 70.94
+ C EU I EN 53.84 80.72 85.89 75.75 60.41 79.20 60.00 79.71 82.40 73.03 81.33 63.60 60.00 73.53 69.62
+ C EU I EU 54.44 82.20 80.33 75.08 59.56 62.00 55.19 80.69 80.21 69.12 75.22 58.00 53.70 72.01 64.73
+ C EU I EN+EU 53.92 81.65 85.78 75.63 60.79 74.40 63.33 80.85 82.06 73.06 80.33 62.80 61.11 73.40 69.41
[…]
+ I EN 53.33 80.68 86.44 74.50 59.22 73.60 61.11 80.09 81.54 72.21 78.11 64.40 58.15 72.73 68.35
+ I EU 54.44 82.41 84.11 74.87 58.63 64.00 62.22 80.09 81.27 70.65 77.22 56.40 54.81 74.19 65.66
+ I EN+EU 53.75 80.89 85.89 74.41 59.14 73.20 63.70 79.22 81.14 72.44 80.44 64.40 56.67 74.39 68.97
+ C EU I EN 54.18 81.94 84.11 74.50 59.48 74.40 59.63 80.14 81.27 71.65 76.33 56.80 57.41 73.33 65.97
+ C EU I EU 53.92 82.53 81.78 74.50 59.09 58.00 62.96 80.09 80.21 69.09 74.89 47.60 55.56 72.60 62.66
+ C EU I EN+EU 54.78 82.24 84.56 74.58 60.62 73.20 61.48 80.63 81.01 71.65 76.78 60.40 59.26 73.20 67.41
[…]
+ I EU 54.18 82.07 90.89 75.38 47.88 80.00 64.81 79.60 80.54 74.09 86.44 70.00 60.37 74.85 72.92
+ I EN+EU 50.85 77.99 91.56 75.59 49.11 81.60 65.19 72.25 69.49 72.10 85.89 73.20 62.96 74.85 74.23
+ C EU 53.84 81.86 90.22 77.05 63.16 68.40 64.81 80.20 83.26 73.74 82.67 61.20 59.63 73.20 69.17
+ C EU I EN 51.96 79.42 91.00 77.38 62.01 81.20 62.59 74.59 76.31 74.35 85.67 73.20 59.26 72.47 72.65
+ C EU I EU 53.58 81.65 90.67 76.55 61.63 72.00 64.07 79.60 80.28 74.65 84.22 70.00 58.52 75.18 71.98
+ C EU I EN+EU 51.62 79.46 90.89 77.13 61.17 84.80 65.93 74.86 74.12 74.65 85.11 72.80 60.00 74.06 72.99
[…]
+ C EU I EN 66.89 87.58 96.33 88.46 74.70 94.80 77.04 81.66 79.02 84.09 93.33 86.00 84.40 81.14 86.82
[…]
D.3 Arena Results
Complete human evaluation results in terms of Bradley–Terry scores and final rankings for each arena dimension (content, language, and global quality) can be consulted in Table 8. Additionally, the detailed global win-rates and battle counts for each model pair are shown in Fig. 12. Note that the reported win-rates incorporate ties in the calculation, resulting in win and loss percentages that do not necessarily sum to 100% for each model pair. Note also that base models (i.e., Base EN and Base EU) were not included in the arena evaluation, as they are not capable of following instructions. However, their performance on benchmarks can be consulted in Section D.2.
colspec=lrrlrlrlrlrl,
cells=font=,
row1,2=font=,
column1,4,8=rightsep=12pt,
column4,8,12=l,font=,leftsep=2pt,preto=(,appto=),
column6,10=l,font=,leftsep=2pt,
rowsep=0pt
[…]
1 1188 +13/-17 1 = 1183 +15/-13 1 = 1093 +12/-10
3.5 Sonnet
2 1153 +13/-21 2 = 1150 +12/-17 3 1 1082 +11/-11
INS EN,70B C EU I EN
3 1141 +15/-11 3 = 1127 +10/-11 2 1 1083 +13/-13
[…]
5 1050 +13/-14 4 1 1047 +12/-12 4 1 1038 +10/-8
[…]
5 1050 +14/-11 5 = 1045 +11/-13 6 1 1034 +8/-10
[…]
6 1038 +13/-13 6 = 1031 +15/-12 5 1 1036 +11/-10
[…]
7 1025 +13/-11 7 = 1026 +13/-13 10 3 1019 +7/-12
[…]
8 1022 +13/-8 8 = 1019 +12/-11 9 1 1021 +10/-10
[…]
9 1017 +12/-11 10 1 1004 +11/-14 7 2 1027 +10/-9
[…]
10 1008 +13/-13 9 1 1008 +10/-13 12 2 1008 +9/-9
[…]
12 1005 +17/-14 13 1 989 +16/-13 8 4 1026 +10/-13
[…]
12 1005 +12/-14 11 1 1000 +13/-13 11 1 1014 +10/-10
[…]
13 991 +13/-16 12 1 990 +13/-17 15 2 991 +11/-10
[…]
14 985 +12/-16 14 = 985 +13/-15 13 1 1002 +10/-10
[…]
15 982 +12/-17 15 = 984 +10/-14 14 1 994 +10/-11
[…]
16 942 +9/-12 17 1 935 +8/-12 16 = 974 +9/-8
[…]
17 936 +14/-11 16 1 943 +10/-13 18 1 944 +11/-10
[…]
18 918 +14/-14 18 = 924 +11/-12 19 1 943 +12/-11
[…]
19 915 +16/-14 19 = 922 +15/-17 17 2 955 +12/-10
[…]
20 896 +14/-16 20 = 909 +12/-16 20 = 925 +10/-13
[…]
21 722 +19/-17 21 = 766 +17/-14 21 = 783 +12/-12
[…]
Appendix E Safety and Bias
We assess the extent to which our instruction-tuning strategy preserves the safety and bias alignment properties of the backbone models. Specifically, we evaluate the model variant INS EN C EU I EN in both 8B and 70B parameter sizes, comparing it with two critical counterparts: (1) BAS EN C EU I EN, to analyze the impact of starting from an already instruction-tune backbone versus a base model (8B only); and (2) Instruct EN, to measure potential alignment changes introduced with our fine-tuning data mix. This evaluation aims to ensure that models maintain appropriate safety guardrails and fairness characteristics in both Basque and English.
colspec=llX[r,2]X[r,1]X[l,1]X[r,1]X[l,1]X[r,2]X[r,1]X[l,1],
colsep=4pt,
cells=font=,
row1,2=font=,
column2,7=rightsep=8pt,
column5,7,10=font=,preto=(,appto=),
cell3-83,8=appto= ,
rowsep=0pt
[…]
Lang Bench
INS EN C EU I EN
[…]
INS EN C EU I EN
[…]
VR ↓ 24.00 44.00 -20.00 8.00 +16.00 4.00 20.00 -16.00
FRR ↓ 0.00 0.00 = 0.00 = 0.00 0.00 =
BBQ ↑ 70.80 51.73 +19.07 71.03 -0.23 85.29 84.90 +0.39
[…]
VR ↓ 16.00 20.00 -4.00 4.00 +12.00 4.00 8.00 -4.00
FRR ↓ 0.00 4.00 -4.00 0.00 = 0.00 0.00 =
BBQ ↑ 87.06 72.35 +14.71 87.65 -0.59 94.38 95.32 +0.96
[…]
Safety.
To test safety, we construct a Basque-language dataset inspired by XSTest Röttger et al. (2024), combining clearly unsafe prompts with superficially similar but safe ones. We measure both Violation rates (VR) and False Refusal rates (FRR) where the model wrongly declines safe prompts. The dataset includes a total of 50 instances, comprising both unsafe and safe prompts across five sensitive categories (self-harm, drugs, child-exploitation, terrorism, explicit-content), adapted to the Basque context and translated into English for cross-language comparison. The outputs of the models were manually annotated by three members performing red teaming. Annotation agreement for unsafe prompt outputs was high (average agreement of 0.973; Fleiss’ Kappa: 0.786; Krippendorff’s Alpha: 0.789). For safe prompt outputs, the annotators unanimously agreed on every item.
Bias.
For bias evaluation, we use BasqBBQ Zulaika and Saralegi (2025) for Basque and BBQ Parrish et al. (2022) for English to analyze disparities across languages. The evaluation is conducted in the same way as described in Section 5.1, using LM Evaluation Harness framework. We use the accuracy metric to evaluate the bias of the model, measuring its ability to choose the correct answer even when biased traps are added to mislead it Parrish et al. (2022).
Evaluation.
As shown in Table 9, the BAS EN C EU I EN model demonstrates low safety and high bias in the Basque language compared to English, reflecting a significant lack of alignment. In contrast, models based on Instruct EN (namely, Instruct EN and INS EN C EU I EN) are better aligned in both languages, as evidenced by the results. Interestingly, in Basque, the 8B Instruct EN shows a lower violation rate than both BAS EN C EU I EN and INS EN C EU I EN models and even the larger 70B counterpart. This behavior is likely due to the model’s limited comprehension of instructions in the Basque language, leading it to produce inconsistent yet safe responses. On the other hand, the 70B models display comparable performance across both languages, with the INS EN C EU I EN version slightly outperforming others and achieving a lower violation rate in Basque. Notably, the false refusal rate remains very low across all models, indicating that safety mechanisms do not come at the cost of excessive conservatism. Bias outcomes in English are consistently better than in Basque, and larger models generally perform better than smaller ones. However, the differences in Bias between the original and language-adapted models are minimal. Overall, we demonstrate that much of the safety and bias alignment is transferred in instruct models, not only in the newly added language, but also in the predominant language.